What is the probability that a randomly chosen female has height between 145 and 157 cm?


1
Expert's answer
2021-02-16T02:58:10-0500

The heights of an adult female population are normally distributed with mean 160 cm and standard deviation 8 cm.

Let "X=" the height of an adult female: "X\sim N(\mu, \sigma^2)."

Then  "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=160\ cm, \sigma=8\ cm."


"P(145<X<157)=P(X<157)-P(X\leq145)"

"=P(Z<\dfrac{157-160}{8})-P(Z\leq\dfrac{145-160}{8})"

"=P(Z<-0.375)-P(Z\leq-1.875)"

"\approx0.35383-0.03040\approx0.3234"

The probability that a randomly chosen female has height between 145 and 157 cm is 0.3234.